OpenStudy (anonymous):

help! Solve the compound inequality 9 + 3n >(underlined) 6 or 5n < 25. n < −1 or n < 5 n > −5 or n > 5 n > −1 or n < 5 n < −5 or n > 5

OpenStudy (anonymous):

by underlined I mean: and also all the answers have the underlined thing too I just could't put it in the question box!

OpenStudy (anonymous):

use ">=" i think most people will know what you mean. 9+3n >= 6 or 5n < 5 9+3n >= 6 -9 3n >= -3 / 3 n >= -1 5n < 5 / 5 n < 5 n >= -1 or n < 5

OpenStudy (anonymous):

oh okay! I never knew another way to write the underlined great/less than symbol. thank you sooo much for you're help!

OpenStudy (anonymous):

you can try \[\leq\]